# Localization of spectral Turán-type theorems

### v1

@@ Abstract
As evidence of their validity, we verify the above conjectures for diamond-free graphs and random graphs. We also propose strengthening of the spectral versions of the Erdős, Stone and Simonovits Theorem by replacing the spectral radius with $\sqrt{s^{+}(G)}$ and establish it for all $F$-free graphs with $\chi(F)=3$. A key ingredient in our proofs is a general upper bound relating $\sqrt{s^{+}(G)}$ to the triangle count $t(G)$. Finally, we prove a localized version of Nikiforov’s walk inequality and conjecture a stronger localized version. These results contribute to the broader program of localizing spectral extremal inequalities.
@@ 1.2 Known results
For any graph $G$, we have
@@
Recently, Li, Liu and Zhang [20] proved an edge-spectral generalization of Erdős-Stone-Simonovits theorem.
@@
For any $F$-free graph $G$ where $\chi(F)\geq 3$. Then
@@
Recently, Liu and Ning generalized Theorem 1.7 to weighted graphs, which implies the results of Bradač, Malec and Tompkins, and Adak and Chandran for Tuŕan’s Theorem.
@@ 1.3 Our contribution
Let $G$ be a graph of order $n$ and clique number $\omega$. Then
$\sqrt{s^{+}(G)}\leq n\left(1-\frac{1}{2\omega-2}\right).$
@@
For any graph $F$-free graph $G$ with chromatic number $\chi(F)\geq 3$, we have
@@
We prove this conjecture for $F$-free graphs with $\chi(F)=3$ (Theorem 3.2).
@@
The above conjecture has been verified for weakly perfect graphs by Ando and Lin [3], triangle-free graphs by Lin, Ning and Wu [21], regular graphs by Zhang [32], and random graphs by Liu and Bu [22]. Motivated by the recent localization results, it is only natural to ask if the following edge-localized refinement of the Bollobás-Nikiforov Conjecture is true.
@@
One can also consider a version of the Bollobás-Nikiforov conjecture for the $F$-free graphs.
@@
For any graph $F$-free graph $G$ with chromatic number $\chi(F)\geq 3$, we have
@@
We prove this statement for $\chi(F)=3$ (Theorem 4.4).
@@ 2. $(ii)$
For any integer $r\geq 1$, we have
$\lambda_{1}^{r}(G)\leq\sum_{v\in V(G)}w_{r}(v)\sqrt{1-\frac{1}{c(v)}}\sqrt{1-\frac{1}{\omega}}.$
@@ 2.1 Weighted Motzkin-Straus type theorems
It is easily seen that the equality case is similar to that of Lemma 2.1. ∎
@@ 2.4 Random graphs
For the random graph $G=G(n,p)$, we have:
@@ 3.1 Some general upper bounds
$\lambda_{1}\leq\sqrt{s^{-}+\frac{6t(G)}{\lambda_{1}}}.$
@@
$\sqrt{s^{+}}=\frac{\lambda_{1}}{\lambda_{1}}\sqrt{s^{+}}\leq\frac{1}{\lambda_{1}}\sqrt{s^{-}+\frac{6t(G)}{\lambda_{1}}}\sqrt{s^{+}}\leq\frac{1}{\lambda_{1}}\frac{2m+\frac{6t(G)}{\lambda_{1}}}{2}=\frac{n}{2}+\frac{3t(G)}{\lambda^{2}_{1}},$
@@
Owing to the difficulty of proving Conjecture 1.1 and 1.2, it is of interest to determine a non-trivial function $f:\mathbb{N}\rightarrow\mathbb{N}$ such that
@@
For Conjecture 1.1, we can improve the function to $f(\omega)=2\omega-2$, which we do below.
@@
Using Lemma 1.10, we see
$\sqrt{s^{+}}\leq\frac{n}{2}+\frac{3t(G)}{\lambda^{2}_{1}}\leq n\left(1-\frac{1}{2\omega-2}\right),$
completing the proof. ∎
@@ 4.1 Diamond-free graphs
If $t(G)=0$, then the assertion is clear. So assume $t(G)\geq 5$.
Since $G$ is diamond-free, for any $e\in E(G)$, $c(e)\in\{2,3\}$. Moreover, if $c(e)=3$, then $e$ is contained in a unique triangle. Thus, we have
@@
Using Theorem 4.2, we have
$\lambda_{1}^{2}+\lambda_{2}^{2}\leq m+\left(\frac{3}{\sqrt{2}}t(G)\right)^{2/3}\leq m+t(G),$
whenever $t(G)\geq 5$. ∎
When $t(G)\in\{1,2,3,4\}$, Theorem 1.10 gives an upper bound which is only slightly worse than $m+t(G)$.

### v2

@@ Abstract
As evidence of their validity, we verify the above conjectures for diamond-free graphs and random graphs. We also propose strengthening of the spectral versions of the Erdős-Stone-Simonovits Theorem by replacing the spectral radius with $\sqrt{s^{+}(G)}$ and establish it for all $F$-free graphs with $\chi(F)=3$. A key ingredient in our proofs is a general upper bound relating $\sqrt{s^{+}(G)}$ to the triangle count $t(G)$. Finally, we prove a localized version of Nikiforov’s walk inequality and conjecture a stronger localized version. These results contribute to the broader program of localizing spectral extremal inequalities.
@@ 1.2 Known results
For any graph $G$ with the clique number $\omega$, we have
@@
Li, Liu and Zhang [21] proved an edge-spectral generalization of Erdős-Stone-Simonovits Theorem.
@@
For any $F$-free graph $G$ where $\chi(F)\geq 3$, we have
@@
Recently, Liu and Ning generalized Theorem 1.7 to weighted graphs, which implies the results of Bradač, and Malec and Tompkins for Tuŕan’s Theorem.
@@ 1.3 Our contribution
Let $G$ be a graph of order $n$ and the clique number $\omega$. Then
$\sqrt{s^{+}(G)}\leq n\sqrt{1-\frac{1}{\omega}-\frac{1}{\omega^{2}}}\ .$
@@
Contrast this with a known result of Elphick and Linz [10] which states that $\sqrt{s^{-}(G)}\leq\frac{n}{2}$ for any graph $G$ of order $n$. This further suggests an asymmetry between $s^{+}$ and $s^{-}$; also refer [37].
@@
For any $F$-free graph $G$ where $\chi(F)\geq 3$, we have
@@
We prove this conjecture for $F$-free graphs when $\chi(F)=3$ (Theorem 3.4).
@@
The above conjecture has been verified for weakly perfect graphs by Ando and Lin [3], triangle-free graphs by Lin, Ning and Wu [24], regular graphs by Zhang [38], and random graphs by Liu and Bu [25]. Kumar and Pragada [20] verified the conjecture for graphs with few triangles, which includes cycle-free graphs, book-free graphs and planar graphs. Zeng and Zhang [36] proved the conjecture for line graphs and also strengthened some results from [20]. Motivated by the recent localization results, it is only natural to ask if the following edge-localized refinement of the Bollobás-Nikiforov Conjecture is true.
@@
We note that equality in Conjecture 1.15 is attained not just by regular complete multipartite graphs (barring $K_{n}$), but also by the disjoint union of two regular complete multipartite graphs of not necessarily the same regularities. This suggests that Conjecture 1.15 captures a rich extremal phenomenon.
Recently, Li, Liu and Zhang [22] proposed a strengthening of the Bollobás-Nikiforov Conjecture for $F$-free graphs where $F$ is a color-critical graph.
Conjecture 1.16 ([22]).
Let $F$ be a color-critical graph with $\chi(F)\geq 4$. Then any $F$-free graph $G$ with sufficiently large size $m$ satisfies
$\lambda_{1}^{2}(G)+\lambda^{2}_{2}(G)\leq 2m\left(1-\frac{1}{\chi(F)-1}\right).$
We propose a weaker bound, but for any graph $F$ with $\chi(F)\geq 3$.
@@
For any $F$-free graph $G$ where $\chi(F)\geq 3$, we have
@@
We prove this statement when $\chi(F)=3$ (Theorem 4.5).
@@ 2. $(ii)$
Let $G$ be a graphFor any integer $r\geq 1$, we have
$\lambda_{1}^{r}(G)\leq\sum_{v\in V(G)}w_{r}(v)\sqrt{1-\frac{1}{c(v)}}\sqrt{1-\frac{1}{\omega(G)}}.$
@@ 2.1 Weighted Motzkin-Straus type theorems
It is easily seen that the equality case is similar to that of Lemma $2.$ ∎
@@ 2.4 Random graphs
For the random graph $G=G(n,p)$, the following hold $a.$s:
@@ 3.1 Some general upper bounds
$\lambda_{1}\leq\sqrt{s^{-}+\frac{6t(G)}{\lambda_{1}}}.$ (1)
@@
$\sqrt{s^{+}}=\frac{\lambda_{1}}{\lambda_{1}}\sqrt{s^{+}}\leq\frac{1}{\lambda_{1}}\sqrt{s^{-}+\frac{6t(G)}{\lambda_{1}}}\sqrt{s^{+}}\leq\frac{1}{\lambda_{1}}\frac{2m+\frac{6t(G)}{\lambda_{1}}}{2}\leq\frac{n}{2}+\frac{3t(G)}{\lambda^{2}_{1}},$
@@
We note here that inequality (1) in the above proof can be rearranged to obtain a lower bound on the number of triangles in a graph:
$t(G)\geq\frac{\lambda_{1}(G)(\lambda_{1}^{2}(G)-s^{-}(G))}{6}.$ (2)
One can compare (2) with the following well-known triangle counting result of Bollobás and Nikiforov [6]: for any graph $G$,
$t(G)\geq\frac{\lambda_{1}(G)(\lambda_{1}^{2}(G)-m)}{3}.$ (3)
Inequality (3) was also independently observed by Cioabă, Feng, Tait and Zhang [8], and the equality case was analyzed by Ning and Zhai [32]; also refer [23].
Owing to the difficulty of proving Conjecture 1.9 and 1.11, it is of interest to determine a non-trivial function $f:\mathbb{N}\rightarrow\mathbb{N}$ such that
@@
For Conjecture 1.9, we can improve the function to $f(\omega)=2\omega-2$, thus proving Theorem 1.10.
@@
Using Theorem 1.12, we see
$\sqrt{s^{+}}\leq\frac{n}{2}+\frac{3t(G)}{\lambda_{1}^{2}}\leq n\left(1-\frac{1}{2\omega-2}.\right)$
Next, we improve the above bound even further. Rearranging inequality 1 and noting that $s^{-}=2m-s^{+}$, we get
$s^{+}\leq 2m-\lambda_{1}^{2}+\frac{6t(G)}{\lambda_{1}}.$
Since $2m\leq n\lambda_{1}$ and substituting the bound for $t(G)$, we have
$s^{+}\leq n\lambda_{1}\left(2-\frac{1}{\omega-1}\right)-\lambda_{1}^{2}$
Wilf’s inequality gives us $\lambda_{1}\leq n\left(1-\frac{1}{\omega}\right)$, so maximing the function on the right gives
$s^{+}\leq n^{2}\left(1-\frac{1}{\omega}-\frac{1}{\omega^{2}}\right).$
This completes the proof. ∎
@@ 4.1 Diamond-free graphs
If $t(G)=0$, then the assertion is known to be true. So assume $t(G)\geq 5$. Since $G$ is diamond-free, for any $e\in E(G)$, $c(e)\in\{2,3\}$. Moreover, if $c(e)=3$, then $e$ is contained in a unique triangle. Thus, we have
@@
Since $t(G)\geq 5$ and using Lemma 4.3, we have
$\lambda_{1}^{2}+\lambda_{2}^{2}\leq m+\left(\frac{3}{\sqrt{2}}t(G)\right)^{2/3}\leq m+t(G).\qed$
When $t(G)\in\{1,2,3,4\}$, Lemma 4.3 gives an upper bound which is only slightly worse than $m+t(G)$.
@@ 5 Walk inequalities
Acknowledgement
M. Rajesh Kannan acknowledges financial support from the ANRF-CRG India and SRC, IIT Hyderabad. The authors thank Clive Elphick for several helpful comments. We also thank Rajat Adak, Yongtao Li and Xiao-Dong Zhang for suggesting some references.